Word problem: At Bake My Day, the owner decorates the day's cake orders three times as fast as her new assistant. If they can decorate all the cake orders working together in 75 min, how long would it take foreach of them to decorate the cakes working individually? Not sure how to set this up.

Answer by stanbon(75887)

Word problem: At Bake My Day, the owner decorates the day's cake orders three times as fast as her new assistant. If they can decorate all the cake orders working together in 75 min, how long would it take foreach of them to decorate the cakes working individually? Not sure how to set this up.
-----
Let the slower baker's time be 3x min/order ; rate = 1/(3x) order/min
The faster baker's time is x min/order ; rate = 1/x order/min
----
The together time is 75 min/order ; rate = 1/75 order/min
---------------------
Equation:
rate + rate = together rate
1/3x + 1/x = 1/75
Multiply thru by 75x to get
25 + 75 = x
x = 100 min (time for the faster baker alone)
3x = 300 min (time for the slower baker alone)
------------------
Cheers,
Stan H.